OpenStudy (anonymous):

Find the perimeter of an isosceles triangle with a vertical angle of 100° and a base of 25cm. Given your answer in two decimal places.

OpenStudy (magbak):

@Katieholmes use this http://www.mathportal.org/calculators.php

OpenStudy (campbell_st):

use the law of sines find the base angles of the isosceles triangles then \[\frac{x}{\sin(40)} = \frac{25}{\sin(100)}\] just solve for x. The perimeter will be P = 2x + 25 hope it helps

OpenStudy (anonymous):

both really helped thank you guys! @campbell_st @magbak

OpenStudy (magbak):

NO BUT @campbell_st is better than me

OpenStudy (anonymous):

Perimeter p = 57.64 http://www.triangle-calculator.com/?what=iso&a=C%3D100+b%3D25&submit=Solve

OpenStudy (anonymous):

thank you too! @triangleguru